# NLPrompt: Noise-Label Prompt Learning for Vision-Language Models

### v2

@@ 3 Preliminary
Notation. In our work, vectors are represented by lowercase bold letters, matrices by uppercase bold letters, and scalars by regular, non-bold letters. The $\ell_{2}$-norm of a vector $\mathbf{v}$ is denoted as $|\mathbf{v}|_{2}$. For matrices, the spectral norm of $\mathbf{A}$ is indicated by $|\mathbf{A}|_{2}$, and the Frobenius norm by $|\mathbf{A}|_{F}$. The indicator function is represented by $\mathds{1}(\cdot)$. Finally, sequences of integers are represented as $[n]=\{1,2,\ldots,n\}$, and sequences of elements, such as vectors, are similarly denoted as $\mathbf{v}_{[n]}=\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{n}\}$.
@@
This problem is a linear programming task, which becomes computationally expensive as the problem scale increases. To address this, Sinkhorn [12] proposed adding an entropic regularization, which allows for a closed-form solution and provides a ”lightspeed” algorithm that only requires iterative scaling of the transportation matrix. The entropic regularized formulation is given as follows:
@@ 4 Theoretical Analysis for the Robustness of PromptMAE
Since the learnable prompt can be expressed as a linear combination of the features, we can analyze the dynamics of these coefficients to understand the learning progress of the prompts. The normalization factor, such as $||\bm{\mu}||_{2}^{-2}$, ensures that the coefficients are comparable to the inner product of the prompt and the features, i.e., $\beta^{(t)}\approx\langle\mathbf{p}^{(t)},\bm{\mu}\rangle$.
@@ 5 Methodology
$\ell_{\text{NLPrompt}}$ $=\sum_{i\in\mathcal{D}_{\text{clean}}}-\mathbf{y}_{i}^{\top}\log\mathbf{s}_{i}+\sum_{j\in\mathcal{D}_{\text{noisy}}}||\mathbf{y}_{j}-\mathbf{s}_{j}||_{1}.$ (11)
@@ 6.2 Noise Setting
For these synthetic noisy datasets, we introduce two types of noise: symmetric noise (denoted as Sym) and asymmetric noise (denoted as Asym). We only flip the labels of the training set in these datasets while keeping the test set unchanged. For symmetric noise, the clean labels in the training set are randomly flipped to other labels with equal probability. This means that labels within the same class can be incorrectly mapped to multiple different classes. For asymmetric noise, the clean labels in the training set are only flipped to a unique neighboring label, with labels within the same class being mapped exclusively to their successor class. Due to its stronger structural nature, asymmetric noise has a more significant negative impact on model performance, making it a stricter robustness test to evaluate the model’s stability and adaptability when confronted with label noise. The goal of learning with noisy labels is to train a robust model on the noisy training set and achieve high accuracy on the clean test set.
@@ C.4 Comparison with JoAPR
The discrepancy in JoAPR results compared to the original paper arises from different noise settings. JoAPR introduces noisy samples uniformly across each class, whereas we randomly distribute noisy samples throughout the entire training set. For example, at a 75% noise rate with 16 shots, JoAPR’s setting turns 12 out of 16 samples in each class into noisy samples and ensuring at least 4 clean samples per class. In contrast, our setting may lead to some classes having only 1 or even no clean samples. Thus, our setting is more challenging and practical, leading to significant performance variance at high noise levels. Additionally, we implemented our NLPrompt on the DTD dataset using JoAPR’s settings and consistently observed improvements over JoAPR. The results are shown in Table A9.
Table A9: Comparison with JoAPR in the JoAPR’s noise setting.
@@ C.6 Computational Overhead of Optimal Transport
We evaluate the computation time of optimal transport through experiments. Leveraging the efficient Sinkhorn algorithm for optimal transport, the computational overhead is minimal. For the Caltech101 dataset under a 16-shot learning setup, we perform batch matching, where 1600 image features are matched against 100 classes. The average time required for optimal transport per epoch is $0.00173$ seconds, while the original backward process takes $4.352$ seconds. Furthermore, even for significantly larger datasets, such as $100,000$ images ×1,000$\times$ 1,000× 1 , 000 classes, the average computation time for the optimal transport method is approximately $1.888$ seconds, which remains relatively small compared to the overall training time.
@@ Appendix E Theoretical Analysis for the Robustness of PromptMAE
Disuccusion on feature decomposition intuition Decomposite the trainable parameters in latent space into linear combinations is a general technique in feature learning theory [1, 3]. Intuitively, in a text prompt, certain ”core” words (such as adjectives describing class features before class names) play a key role in determining the image classification and are considered ”task-relevant,” while other words are ”task-irrelevant.”
@@
$\ell_{\text{MAE}}(\mathbf{s}_{i},\mathbf{y}_{i})=\sum\limits_{c=1}^{C}|y_{i,c}-s_{i,c}|=||\mathbf{y_{i}-\mathbf{s_{i}}}||_{1}$ (13)
@@
  • The learning rate $\eta\leq\tilde{O}(\min\{||\bm{\mu}||_{2}^{2},\sigma_{p}^{-2}m^{-1}\})$ and the standard deviation of network weight initialization $\sigma_{0}\leq\tilde{O}(mn)\cdot\min\{(||\bm{\mu}||_{2}^{2},\sigma_{p}\sqrt{d}^{-1})\}$.
@@
Gradient analysis. In the definition of the text encoder (3), the incorporation of $\mathbf{W}\mathbf{p}_{c}$ introduces nonlinearity between the trainable prompt and the class prompt while maintaining the overall function’s nonlinear nature. The assumptions regarding the image encoder (5) suggest that task-relevant features differ depending on whether the label is positive or negative, while task-irrelevant features remain arbitrary and independent of the label’s polarity. Furthermore, the training loss objective is designed to strengthen the similarity between the image feature $g(\mathbf{x}_{i})$ and the text feature generated by the label class prompt $\mathbf{p}_{y_{i}}$. This section discusses the computational approach used to analyze the performance of the algorithm, focusing on the gradient calculations essential for optimizing the model parameters. Leveraging the properties of the gradient of the Softmax function, we have:
@@
$\beta^{(t+1)}$ $=\beta^{(t)}+\frac{\eta}{n}\sum\limits_{i=1}^{n}\ell^{\prime}_{i}\sigma^{\prime}_{1,i}\tilde{y}_{i}y_{i}||\bm{\mu}||_{2}^{2}$ (24)
$=\beta^{(t)}+\frac{\eta}{n}\sum\limits_{i\in S_{+}}\ell^{\prime}_{i}\sigma^{\prime}_{1,i}||\bm{\mu}||_{2}^{2}-\frac{\eta}{n}\sum\limits_{i\in S_{-}}\ell^{\prime}_{i}\sigma^{\prime}_{1,i}||\bm{\mu}||_{2}^{2}$
$\phi_{l}^{(t+1)}$ $=\phi^{(t)}+\frac{\eta}{n}\sum\limits_{i=1}^{n}\ell^{\prime}_{i}\sigma^{\prime}_{1,i}\tilde{y}_{i}x_{l+1,i}||\bm{\xi}_{l}||_{2}^{2}.$ (25)
@@ E.1 Theoretical analysis
$\mathbf{W}\mathbf{p}=\begin{bmatrix}\bm{\mu}^{\top}\mathbf{p}\\ \bm{\xi}_{1}^{\top}\mathbf{p}\\ \vdots\\ \bm{\xi}_{L}^{\top}\mathbf{p}\end{bmatrix}=\begin{bmatrix}\beta||\bm{\mu}||_{2}\\ \phi_{1}||\bm{\xi}_{1}||_{2}\\ \vdots\\ \phi_{L}||\bm{\xi}_{L}||_{2}\end{bmatrix}.$ (30)
Also, since the weight and class prompts are fixed during prompt learning, $\mathbf{Wp}_{+}$ and $\mathbf{Wp}_{-}$ can be treated as two constant terms. To assess the algorithm’s performance, we evaluate the error rate during the testing procedure, which serves as our test loss, denoted as $\ell_{\mathcal{D}}$:
@@
$\beta^{(t+1)}_{\text{CE}}$ $=\beta^{(t)}_{\text{CE}}+\eta\left[(1-p)\frac{1}{\mathds{E}[s_{y}]}||\bm{\mu}||_{2}^{2}-p\frac{1}{1-\mathds{E}[s_{y}]}||\bm{\mu}||_{2}^{2}\right],$ (37)
@@
$\beta^{(t+1)}_{\text{MAE}}$ $=\beta^{(t)}_{\text{MAE}}+\eta\left[(1-p)\cdot 2\cdot||\bm{\mu}||_{2}^{2}-p\cdot 2\cdot||\bm{\mu}||_{2}^{2}\right],$ (39)
@@
Given that $\mathds{E}[s_{y}]>\frac{1}{2}$, we have $\frac{1}{1-\mathds{E}[s_{y}]}>2$ and $02\mathds{E}[s_{y}]-1>$ 0. Additionally, since $0\leq p\leq\frac{1}{2}$, it follows that:
@@
$\frac{\beta^{(t+1)}_{\text{MAE}}}{\phi^{(t+1)}_{\text{MAE}}}=\frac{\beta^{(t)}_{\text{MAE}}+\Delta\beta_{\text{MAE}}^{(t)}}{\phi^{(t)}_{\text{MAE}}+\Delta\phi_{\text{MAE}}^{(t)}}\geq\frac{\beta^{(t)}_{\text{MAE}}}{\phi_{\text{MAE}}^{(t)}}+\frac{\Delta\beta^{(t)}_{\text{MAE}}}{\phi_{\text{MAE}}^{(t)}}\geq\frac{\beta^{(t)}_{\text{CE}}}{\phi_{\text{CE}}^{(t)}}+\frac{\Delta\beta^{(t)}_{\text{CE}}}{2\mathds{E}[s_{y}]\phi_{\text{MAE}}^{(t)}}\geq\frac{\beta^{(t)}_{\text{CE}}}{\phi_{\text{CE}}^{(t)}}+\frac{\Delta\beta^{(t)}_{\text{CE}}}{\phi_{\text{CE}}^{(t)}}\geq\frac{\beta^{(t+1)}_{\text{CE}}}{\phi^{(t+1)}_{\text{CE}}}.$ (48)
Additionally, we have:

### v3

@@ 3 Preliminary
Notation. In our work, vectors are represented by lowercase bold letters, matrices by uppercase bold letters, and scalars by regular, non-bold letters. The $\ell_{2}$-norm of a vector $\mathbf{v}$ is denoted as $|\mathbf{v}|_{2}$. For matrices, the spectral norm of $\mathbf{A}$ is indicated by $|\mathbf{A}|_{2}$, and the Frobenius norm by $|\mathbf{A}|_{F}$. The indicator function is represented by $\mathds{1}(\cdot)$. Finally, sequences of integers are represented as $n]=\{1,2,\ldots,n\}$, and sequences of elements, such as vectors, are similarly denoted as $\mathbf{v}_{[n]}=\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{n}\}$.
@@
This problem is a linear programming task, which becomes computationally expensive as the problem scale increases. To address this, Sinkhorn [12] proposed adding an entropic regularization, which allows for a closed-form solution and provides a "lightspeed" algorithm that only requires iterative scaling of the transportation matrix. The entropic regularized formulation is given as follows:
@@ 4 Theoretical Analysis for the Robustness of PromptMAE
Since the learnable prompt can be expressed as a linear combination of the features, we can analyze the dynamics of these coefficients to understand the learning progress of the prompts. The normalization factor, such as $\bm{\mu}||_{2}^{-2}$, ensures that the coefficients are comparable to the inner product of the prompt and the features, i.e., $\beta^{(t)}\approx\langle\mathbf{p}^{(t)},\bm{\mu}\rangle$.
@@ 5 Methodology
[…]
@@ 6.2 Noise Setting
For these synthetic noisy datasets, we introduce two types of noise: symmetric noise (denoted as Sym) and asymmetric noise (denoted as Asym). We only flip the labels of the training set in these datasets while keeping the test set unchanged. For symmetric noise, the clean labels in the training set are randomly flipped to other labels with equal probability. This means that labels within the same class can be incorrectly mapped to multiple different classes. For asymmetric noise, the clean labels in the training set are only flipped to a unique neighboring label, with labels within the same class being mapped exclusively to their successor class. Due to its stronger structural nature, asymmetric noise has a more significant negative impact on model performance, making it a stricter robustness test to evaluate the model's stability and adaptability when confronted with label noise. The goal of learning with noisy labels is to train a robust model on the noisy training set and achieve high accuracy on the clean test set.
@@ C.4 Comparison with JoAPR
The discrepancy in JoAPR results compared to the original paper arises from different noise settings. JoAPR introduces noisy samples uniformly across each class, whereas we randomly distribute noisy samples throughout the entire training set. For example, at a 75% noise rate with 16 shots, JoAPR's setting turns 12 out of 16 samples in each class into noisy samples and ensuring at least 4 clean samples per class. In contrast, our setting may lead to some classes having only 1 or even no clean samples. Thus, our setting is more challenging and practical, leading to significant performance variance at high noise levels. Additionally, we implemented our NLPrompt on the DTD dataset using JoAPR’s settings and consistently observed improvements over JoAPR. The results are shown in Table A9.
Table A9: Comparison with JoAPR in the JoAPR's noise setting.
@@ C.6 Computational Overhead of Optimal Transport
We evaluate the computation time of optimal transport through experiments. Leveraging the efficient Sinkhorn algorithm for optimal transport, the computational overhead is minimal. For the Caltech101 dataset under a 16-shot learning setup, we perform batch matching, where 1600 image features are matched against 100 classes. The average time required for optimal transport per epoch is $0.00173$ seconds, while the original backward process takes $4.352$ seconds. Furthermore, even for significantly larger datasets, such as $100,000$ images $\times 1,000$ classes, the average computation time for the optimal transport method is approximately $1.888$ seconds, which remains relatively small compared to the overall training time.
@@ Appendix E Theoretical Analysis for the Robustness of PromptMAE
Disuccusion on feature decomposition intuition Decomposite the trainable parameters in latent space into linear combinations is a general technique in feature learning theory [1, 3]. Intuitively, in a text prompt, certain "core" words (such as adjectives describing class features before class names) play a key role in determining the image classification and are considered "task-relevant," while other words are "task-irrelevant."
@@
[…]
@@
[…]
@@
Gradient analysis. In the definition of the text encoder (3), the incorporation of $\mathbf{W}\mathbf{p}_{c}$ introduces nonlinearity between the trainable prompt and the class prompt while maintaining the overall function's nonlinear nature. The assumptions regarding the image encoder (5) suggest that task-relevant features differ depending on whether the label is positive or negative, while task-irrelevant features remain arbitrary and independent of the label's polarity. Furthermore, the training loss objective is designed to strengthen the similarity between the image feature $g(\mathbf{x}_{i})$ and the text feature generated by the label class prompt $\mathbf{p}_{y_{i}}$. This section discusses the computational approach used to analyze the performance of the algorithm, focusing on the gradient calculations essential for optimizing the model parameters. Leveraging the properties of the gradient of the Softmax function, we have:
@@
[…]
@@ E.1 Theoretical analysis
[…]
Also, since the weight and class prompts are fixed during prompt learning, $\mathbf{Wp}_{+}$ and $\mathbf{Wp}_{-}$ can be treated as two constant terms. To assess the algorithm's performance, we evaluate the error rate during the testing procedure, which serves as our test loss, denoted as $\ell_{\mathcal{D}}$:
@@
[…]
@@
[…]
@@
Given that $\mathds{E}[s_{y}]>\frac{1}{2}$, we have $\frac{1}{1-\mathds{E}[s_{y}]}>2$ and $2\mathds{E}[s_{y}]-1>0$. Additionally, since $0\leq p\leq\frac{1}{2}$, it follows that:
@@
$\frac{\beta^{(t+1)}_{\text{MAE}}}{\phi^{(t+1)}_{\text{MAE}}}=\frac{\beta^{(t)}_{\text{MAE}}+\Delta\beta_{\text{MAE}}^{(t)}}{\phi^{(t)}_{\text{MAE}}+\Delta\phi_{\text{MAE}}^{(t)}}\geq\frac{\beta^{(t)}_{\text{CE}}+\Delta\beta^{(t)}_{\text{CE}}}{\phi_{\text{CE}}^{(t)}+\Delta\phi_{\text{CE}}^{(t)}}=\frac{\beta^{(t+1)}_{\text{CE}}}{\phi^{(t+1)}_{\text{CE}}}$. (48)
This inequality arises from the fact that $\frac{\beta_{\text{MAE}}^{(t)}}{\phi_{\text{MAE}}^{(t)}}\geq\frac{\beta_{\text{CE}}^{(t)}}{\phi_{\text{CE}}^{(t)}}$ and the increments satisfy $\frac{\Delta\beta_{\text{MAE}}^{(t)}}{\Delta\phi_{\text{MAE}}^{(t)}}\geq\frac{\Delta\beta_{\text{CE}}^{(t)}}{\Delta\phi_{\text{CE}}^{(t)}}$. Additionally, we have: